# Bayesian Optimization with Gaussian Processes to Accelerate Stationary Point Searches

### 16 Mar 2026 v2

@@ Abstract
Accelerating the explorations of stationary points on potential energy surfaces building local surrogates spans decades of effort. Done correctly, surrogates reduce required evaluations by an order of magnitude while preserving the accuracy of the underlying theory. We present a unified Bayesian Optimization view of minimization, single point saddle searches, and double ended saddle searches through a unified six-step surrogate loop, differing only in the inner optimization target and acquisition criterion. The framework uses Gaussian process regression with derivative observations, inverse-distance kernels, and active learning. The Optimal Transport GP extensions of farthest point sampling with Earth mover’s distance, MAP regularization via variance barrier and oscillation detection, and adaptive trust radius form concrete extensions of the same basic methodology, improving accuracy and efficiency. We also demonstrate random Fourier features decouple hyperparameter training from predictions enabling favorable scaling for high-dimensional systems. Accompanying pedagogical Rust code demonstrates that all applications use the exact same Bayesian optimization loop, bridging the gap between theoretical formulation and practical execution.
@@ 1 Introduction
The rate depends exponentially on the barrier $E_{\text{SP}}-E_{\text{min}}$, with the vibrational frequency ratio ($\nu_{i}^{\text{min}}$ at the minimum, $\nu_{i}^{\text{SP}}$ at the saddle) acting as a prefactor that captures the width of the two basins. The denominator runs over 3N13N-1 modes because the saddle has one imaginary frequency along the reaction coordinate under four assumptions covered elsewhere [82]. In practice, predicting the rate reduces to finding the minima and saddle points 111along with the Hessian test for the first order saddle point estimate, often skipped in high throughput workflows. To this end we consider minimization on an energy surface, along with two common modalities to find saddle points. To this end we consider minimization on an energy surface, along with two common modalities to find saddle points. We will use the term “point searches” to cover the union of the search methods.
The global minimum for such systems being largely ill defined due to the lack of an absolute zero, so the map from geometry to energy and atomistic forces (the pointwise derivative w.r.t positions) also ends up being relative. One can construct several alternative surrogate energy surfaces that perfectly preserve the minima and saddles of the true PES, even if those surrogates distort the broader configurational space. This inherently local understanding of the process provides the leeway required to handle the extreme dimensionality of the $3N$ space despite there being “no-free-lunch” [48] and has connections to vibrational analysis [63].
Point searches typically need hundreds of such evaluations to converge, and this cost becomes prohibitive in large-scale studies where workflow-driven screening with tools such as AiiDA [77] or Snakemake [68] may require characterizing thousands of distinct transitions. may require characterizing thousands of distinct transitions. The problem is compounded in applications that embed saddle searches as an inner loop. Adaptive kinetic Monte Carlo (AKMC) [41, 32] discovers escape routes on the fly from each visited minimum, reaction network exploration catalogs competing pathways in complex catalytic cycles, and high-throughput screening for materials design may require characterizing saddle points across hundreds of candidate systems. In each case, the per-search cost of electronic structure evaluations is the rate-limiting step, and reducing it from hundreds of calls to tens opens qualitatively different scales of investigation.
@@
The second strategy, and the subject of this review, constructs a local, ephemeral surrogate of the PES on-the-fly during each individual search, using only the data generated in the course of that specific calculation. This is an active learning approach [44, 96] in which the surrogate model decides where to sample next, balancing exploitation of the current prediction against exploration of uncertain regions [93, 96]. The surrogate need not represent the PES globally; it only needs to be accurate near the path being optimized. Convergence to a saddle point typically requires on the order of 30 electronic structure evaluations [30], compared to the thousands needed for even a modest MLIP. The surrogate is discarded after each search completes.
@@
Gaussian process regression (GPR) [27, 35] is well matched to this local surrogate role. The GP posterior provides two quantities required by the active learning loop: the predicted energy surface (the posterior mean) and a calibrated uncertainty estimate (the posterior variance). The variance enables active learning. Many optimization steps can be run on the cheap surrogate surface, and when the next expensive electronic structure evaluation is needed, the configuration where the GP is least certain is selected. This active learning loop (evaluate, update the model, optimize on the surrogate, evaluate again where uncertainty is highest) distinguishes the approach from passive regression on a pre-existing database.
@@
The two approaches differ in what the GP is asked to do. In the MLIP setting, the GP is a regression tool that, given a large, pre-computed training set, interpolates to approximate the energy at new configurations. In the active learning setting developed here, the GP is an acquisition tool that decides where to acquire new data based on its own uncertainty, and the training set is assembled incrementally as a byproduct of the search itself. The kernel operates on inverse interatomic distances rather than high-dimensional structural descriptors like the smoothed overlap of atomic positions, or SOAP [10], because the model only needs to resolve the PES in a local neighborhood where pairwise distance features suffice. The inverse-distance kernel has been validated across 500+ reaction benchmarks with both SE and Matern kernels [30, 53]. The computational overhead of the GP (dominated by the $\mathcal{O}(M^{3})$ Cholesky decomposition with $M\sim 30$) is negligible compared to the electronic structure cost it replaces, whereas an MLIP-scale GP with $M\sim 10^{4}$ would require sparse approximations.
@@
This tutorial review provides a self-contained treatment of GPR-accelerated stationary point searches. We develop the mathematical foundations of GPR with derivative observations in sufficient detail for a practitioner to implement the method from scratch, and apply the framework to the dimer method, the nudged elastic band (NEB), and local minimization. Accompanying Rust code (chemgp-core, available at https://github.com/HaoZeke/ChemGP) is both the pedagogical reference and the production implementation; each equation maps to a specific function, and the same binary runs the benchmarks reported here. We give practical guidance on hyperparameter selection, coordinate systems, trust regions, data management, and the statistical evaluation of algorithmic performance. A unifying observation runs through the three applications: every GP method shares the same Bayesian optimization outer loop of training a surrogate, selecting a query point by an acquisition criterion, evaluating the oracle. With four shared components (FPS subset, EMD trust, RFF approximation, LCB acquisition) that are formalized as the Bayesian surrogate loop in Section 4 and developed in detail within the OT-GP framework (Section 8). The unified Bayesian optimization framework using Gaussian process surrogates is detailed in Alg. 1.
@@
Input: Initial configuration(s) $\mathbf{x}_{0}$, oracle tolerance $\epsilon$, trust radius $\Delta$
Output: Converged stationary point(s) $\mathbf{x}^{*}$
1 Initialize dataset $\mathcal{D}\leftarrow\{(\mathbf{x}_{0},V(\mathbf{x}_{0}),\nabla V(\mathbf{x}_{0}))\}$;
2 $\mathbf{x}^{*}\leftarrow\mathbf{x}_{0}$;
3 while $|\nabla V(\mathbf{x}^{*})|>\epsilon$ do
4 Select training subset $\mathcal{D}_{\text{train}}\subset\mathcal{D}$ via farthest point sampling;
5 Train hyperparameters $\boldsymbol{\theta}$ via MAP estimation;
6 Build surrogate model $V_{\text{GP}}$ (Exact GP or Random Fourier Features);
7 Optimize $\mathbf{x}_{\text{prop}}$ on $V_{\text{GP}}$ via method-specific inner loop;
8 Clip $\mathbf{x}_{\text{prop}}$ to trust region boundary $\Delta$ (EMD or Euclidean);
9 Acquire new oracle evaluation at $\mathbf{x}_{\text{prop}}$ based on acquisition criterion;
10   Update dataset $\mathcal{D}\leftarrow\mathcal{D}\cup\{(\mathbf{x}_{\text{prop}},V(\mathbf{x}_{\text{prop}}),\nabla V(\mathbf{x}_{\text{prop}}))\}$;
11   Update trust radius $\Delta$ based on surrogate accuracy;
12 $\mathbf{x}^{*}\leftarrow\mathbf{x}_{\text{prop}}$;
13
14 end while
return Converged stationary point(s) $\mathbf{x}^{*}$
@@ 12 $\mathbf{x}^{*}\leftarrow\mathbf{x}_{\text{prop}}$;
The review is organized as follows. Section 2 establishes the physical setting and the classical search algorithms. Section 3 develops the GPR framework, including molecular kernels and gradient observations. Section 4 formalizes the Bayesian surrogate loop that unifies the three applications. Section 5 presents GPR-accelerated minimum mode following (the GP-dimer), Section 6 covers GPR-accelerated NEB, and Section 7 treats GPR-accelerated minimization. Section 8 introduces the Optimal Transport GP (OT-GP) extensions that address scaling and stability. Section 9 connects the mathematics to executable code and describes the analysis tooling.
@@ 2.3.2 Translation
Input: Midpoint $\mathbf{R}$, initial dimer axis $\hat{\mathbf{N}}$, separation $\Delta R$, force threshold $\epsilon_{\text{force}}$
Output: saddle point at $\mathbf{R}$
1 repeat
2   Place endpoints $\mathbf{R}_{1,2}=\mathbf{R}\pm\Delta R\,\hat{\mathbf{N}}$;
3   Evaluate $\mathbf{F}(\mathbf{R}_{1}),\mathbf{F}(\mathbf{R}_{2})$ on true PES;
 // Rotation
4 repeat
5    $C(\hat{\mathbf{N}})\leftarrow(\mathbf{F}_{2}-\mathbf{F}_{1})\cdot\hat{\mathbf{N}}/\Delta R$;
6      Rotate $\hat{\mathbf{N}}$ to minimize $C$ (CG, Eq. 2.7);
7      Evaluate $\mathbf{F}(\mathbf{R}_{1})$ at new endpoint;
8
9 until rotation converged;
 // Translation
10 if $C(\hat{\mathbf{N}})<0$ then
11    $\mathbf{F}^{\dagger}\leftarrow\mathbf{F}(\mathbf{R})-2[\mathbf{F}(\mathbf{R})\cdot\hat{\mathbf{N}}]\hat{\mathbf{N}}$;
12
13 else
14    $\mathbf{F}^{\dagger}\leftarrow-[\mathbf{F}(\mathbf{R})\cdot\hat{\mathbf{N}}]\hat{\mathbf{N}}$;
15
16   end if
17 $\mathbf{R}\leftarrow\mathbf{R}+\text{L-BFGS step on }\mathbf{F}^{\dagger}$;
18
19until $|\mathbf{F}(\mathbf{R})|<\epsilon_{\text{force}}$;
@@ 2.4.1 The Nudged Elastic Band
A naive elastic band, where springs connect the images and the full potential force acts on each one, fails to converge to the MEP for two reasons. First, the potential force has a component along the path that drags images away from the saddle region and bunches them in low-energy basins (the “sliding-down” artifact). Second, the spring force has a component perpendicular to the path that pulls the chain off the MEP into straight-line shortcuts through high-energy regions (the “corner-cutting” artifact) [91]. The NEB eliminates both by projecting each type of force onto the subspace where it belongs. The true force acts only perpendicular to the path, driving each image toward the MEP, while the spring force acts only parallel to the path, controlling image spacing. The total NEB force on image $i$ is:
@@
This decoupling of shape optimization (perpendicular) from spacing control (parallel) is the “nudging” that gives the method its name.
@@
When the energy is monotonically increasing or decreasing through image $i$, the tangent points toward the uphill neighbor. When image $i$ sits at a local extremum (the “otherwise” case), an energy-weighted average smoothly interpolates between the two directions. The stability condition for the NEB requires that the parallel spring force be bounded by the product of the perpendicular curvature and the image spacing; the bisection tangent violates this at large $P$, while the improved tangent does not.
@@ 2.4.2 The Climbing Image and Its Connection to Minimum Mode Following
Input: Initial chain $\{\mathbf{R}_{0},\ldots,\mathbf{R}_{P}\}$, spring constant $k$, CI threshold $\epsilon_{\text{CI}}$, convergence tolerance $\epsilon_{\text{tol}}$
Output: converged MEP $\{\mathbf{R}_{0},\ldots,\mathbf{R}_{P}\}$
1 CI $\leftarrow$ false;
2 repeat
3   Evaluate $V(\mathbf{R}_{i}),\mathbf{F}(\mathbf{R}_{i})$ at all movable images;
4 for each image $i=1,\ldots,P-1$ do
5      Compute tangent $\hat{\boldsymbol{\tau}}_{i}$ (Eq. 2.13);
6    if CI is true and $i=i_{\max}$ then
7       $\mathbf{F}_{i}\leftarrow-\nabla V+2(\nabla V\cdot\hat{\boldsymbol{\tau}}_{i})\hat{\boldsymbol{\tau}}_{i}$;
8
9    else
10       $\mathbf{F}_{i}^{\perp}\leftarrow-\nabla V+(\nabla V\cdot\hat{\boldsymbol{\tau}}_{i})\hat{\boldsymbol{\tau}}_{i}$;
11       $F_{i}^{s}\leftarrow k(|\mathbf{R}_{i+1}-\mathbf{R}_{i}|-|\mathbf{R}_{i}-\mathbf{R}_{i-1}|)$;
12       $\mathbf{F}_{i}\leftarrow\mathbf{F}_{i}^{\perp}+F_{i}^{s}\,\hat{\boldsymbol{\tau}}_{i}$;
13
14      end if
15
16   end for
17  Update all images $\mathbf{R}_{i}\leftarrow\mathbf{R}_{i}+\Delta t\,\mathbf{F}_{i}$;
18 if CI is false and $\max_{i}\|\mathbf{F}_{i}\|<\epsilon_{\text{CI}}$ then
19      CI $\leftarrow$ false;
20    $i_{\max}\leftarrow\arg\max_{i}V(\mathbf{R}_{i})$;
21
22   end if
23
24until $\max_{i}\|\mathbf{F}_{i}\|<\epsilon_{\text{tol}}$;

### 16 Apr 2026 v3

@@ Abstract
Building local surrogates to accelerate stationary point searches on potential energy surfaces spans decades of effort. Done correctly, surrogates can reduce the number of expensive electronic structure evaluations by roughly an order of magnitude while preserving the accuracy of the underlying theory; the gain depends on oracle cost, search distance, and the availability of analytical forces, and is largest for saddle searches and smallest for minimization near a quadratic basin. We present a unified Bayesian optimization view of minimization, single-point saddle searches, and double-ended path searches: all three share one six-step surrogate loop and differ only in the inner optimization target and the acquisition criterion. The framework uses Gaussian process regression with derivative observations, inverse-distance kernels, and active learning, and we develop optional extensions for production use, including farthest-point sampling with the Earth Mover’s Distance, MAP regularization, an adaptive trust radius, and random Fourier features for scaling. Accompanying pedagogical Rust code demonstrates that all three applications use the same Bayesian optimization loop, bridging the gap between theoretical formulation and practical execution.
@@ 1 Introduction
The rate depends exponentially on the barrier $E_{\text{SP}}-E_{\text{min}}$, with the vibrational frequency ratio ($\nu_{i}^{\text{min}}$ at the minimum, $\nu_{i}^{\text{SP}}$ at the saddle) acting as a prefactor that captures the width of the two basins. The denominator runs over 3N13N-1 modes because the saddle has one imaginary frequency along the reaction coordinate under four assumptions covered elsewhere [84]. In practice, predicting the rate reduces to finding the minima and saddle points 111along with the Hessian test for the first order saddle point estimate, often skipped in high throughput workflows. To this end we consider minimization on an energy surface, along with two common modalities to find saddle points. We will use the term "point searches" to cover the union of the search methods.
The global minimum for such systems being largely ill defined due to the lack of an absolute zero, so the map from geometry to energy and atomistic forces (the pointwise derivative w.r.t positions) also ends up being relative. One can construct several alternative surrogate energy surfaces that perfectly preserve the minima and saddles of the true PES, even if those surrogates distort the broader configurational space. This inherently local understanding of the process provides the leeway required to handle the extreme dimensionality of the $3N$ space despite there being "no-free-lunch" [49] and has connections to vibrational analysis [64].
Point searches typically need hundreds of such evaluations to converge, and this cost becomes prohibitive in large-scale studies where workflow-driven screening with tools such as AiiDA [78] or Snakemake [69] may require characterizing thousands of distinct transitions. The problem is compounded in applications that embed saddle searches as an inner loop. Adaptive kinetic Monte Carlo (AKMC) [42, 33] discovers escape routes on the fly from each visited minimum, reaction network exploration catalogs competing pathways in complex catalytic cycles, and high-throughput screening for materials design may require characterizing saddle points across hundreds of candidate systems. In each case, the per-search cost of electronic structure evaluations is the rate-limiting step, and reducing it from hundreds of calls to tens opens qualitatively different scales of investigation.
@@
The second strategy, and the subject of this review, constructs a local, ephemeral surrogate of the PES on-the-fly during each individual search, using only the data generated in the course of that specific calculation. This is an active learning approach [45, 98] in which the surrogate model decides where to sample next, balancing exploitation of the current prediction against exploration of uncertain regions [95, 98]. The surrogate need not represent the PES globally; it only needs to be accurate near the path being optimized. Convergence to a saddle point typically requires on the order of 30 electronic structure evaluations [31, 30], compared to the thousands needed for even a modest MLIP. The surrogate is discarded after each search completes.
Several other groups have built per-search GP surrogates along similar lines: FLARE and committee-based GAP for on-the-fly molecular dynamics [100, 4], restricted-variance and gradient-enhanced kriging for geometry optimization [81, 24], and earlier GPR-accelerated NEB and dimer codes [56, 54, 15, 102, 27]. What this tutorial adds is the unification: minimization, dimer, and NEB share one Bayesian optimization loop, and the differences reduce to the inner optimizer and the acquisition rule.
@@
Gaussian process regression (GPR) [28, 36] is well matched to this local surrogate role. The GP posterior provides two quantities used by the outer loop: the predicted energy surface (the posterior mean, which plays the role of a cheap surrogate PES on which standard optimizers can run) and a posterior variance (which tracks distance from the training data in the geometry induced by the kernel). The variance is not a direct measure of accuracy against the true PES; it is a self-consistent statement of the surrogate’s own disagreement with itself under its fitted hyperparameters, and it is driven down by sampling more (Section~3). We use it where this is what we need, namely to detect that a region is under-sampled (the LCB stopping rule in the inner loop) and, in NEB, to prioritize image selection when several candidates are available; we do not use it as a per-proposal gate on the oracle, and we pair it with a geometric trust radius that encodes physical displacement bounds directly (Section~7).
@@
The two approaches differ in what the GP is asked to do. In the MLIP setting, the GP is a regression tool that, given a large, pre-computed training set, interpolates to approximate the energy at new configurations. In the active learning setting developed here, the GP is a local surrogate that is optimized on directly, with a trust region bounding each step and an oracle call concluding every outer iteration; the training set is assembled incrementally as a byproduct of the search itself. The kernel operates on inverse interatomic distances rather than high-dimensional structural descriptors like the smoothed overlap of atomic positions, or SOAP [11], because the model only needs to resolve the PES in a local neighborhood where pairwise distance features suffice. The inverse-distance kernel has been validated across 500+ reaction benchmarks with both SE and Matern kernels [31, 30, 54]. The computational overhead of the GP (dominated by the $\mathcal{O}(M^{3})$ Cholesky decomposition with $M\sim 30$) is negligible compared to the electronic structure cost it replaces, whereas an MLIP-scale GP with $M\sim 10^{4}$ would require sparse approximations.
@@
This tutorial review provides a self-contained treatment of GPR-accelerated stationary point searches. We develop the mathematical foundations of GPR with derivative observations in sufficient detail for a practitioner to implement the method from scratch, and apply the framework to the dimer method, the nudged elastic band (NEB), and local minimization. Accompanying Rust code (chemgp-core, source at https://github.com/lode-org/ChemGP and documentation at https://lode-org.github.io/ChemGP/) is the pedagogical reference implementation of every algorithm in the review; each equation maps to a specific function, and the crate runs the illustrative benchmarks reported here. Production-scale saddle-point studies on hundreds of molecular reactions use the C++ gpr_optim code, reported in detail elsewhere [31, 30], and the two implementations share the same algorithmic core. We give practical guidance on hyperparameter selection, coordinate systems, trust regions, data management, and the statistical evaluation of algorithmic performance. A unifying observation runs through the three applications: every GP method shares the same Bayesian optimization outer loop of training a surrogate, selecting a query point by an acquisition criterion, evaluating the oracle. With four shared components (FPS subset, EMD trust, RFF approximation, LCB acquisition) that are formalized as the Bayesian surrogate loop in Section 4 and developed in detail within the OT-GP framework (Section 8). Algorithm 1 gives the high-level skeleton shared by minimization, the dimer, and NEB; the detailed unified framework with its acquisition, training, and trust-region steps appears as Algorithm 4 in Section 4.
@@
1:Initial configuration(s) $\mathbf{x}_{0}$, oracle tolerance $\epsilon$, trust radius $\Delta$
2:Initialize dataset $\mathcal{D}\leftarrow\{(\mathbf{x}_{0},V(\mathbf{x}_{0}),\nabla V(\mathbf{x}_{0}))\}$
[…]
4:while $|\nabla V(\mathbf{x}^{*})|>\epsilon$ do
5:  Select training subset $\mathcal{D}_{\text{train}}\subset\mathcal{D}$ via farthest point sampling
6:  Train hyperparameters $\boldsymbol{\theta}$ via MAP estimation
7:  Build surrogate model $V_{\text{GP}}$ (Exact GP or Random Fourier Features)
8:  Optimize $\mathbf{x}_{\text{prop}}$ on $V_{\text{GP}}$ via method-specific inner loop
9:  Clip $\mathbf{x}_{\text{prop}}$ to trust region boundary $\Delta$ (EMD or Euclidean)
10:  Acquire new oracle evaluation at $\mathbf{x}_{\text{prop}}$ based on acquisition criterion
11:  Update dataset $\mathcal{D}\leftarrow\mathcal{D}\cup\{(\mathbf{x}_{\text{prop}},V(\mathbf{x}_{\text{prop}}),\nabla V(\mathbf{x}_{\text{prop}}))\}$
12:  Update trust radius $\Delta$ based on surrogate accuracy
13:  $\mathbf{x}^{*}\leftarrow\mathbf{x}_{\text{prop}}$
14:end while
15:return Converged stationary point(s) $\mathbf{x}^{*}$
The review is organized as follows. Section 2 establishes the physical setting and the classical search algorithms. Section 3 develops the GPR framework, including molecular kernels and gradient observations. Section 4 formalizes the Bayesian surrogate loop that unifies the three applications. Section 5 presents GPR-accelerated minimum mode following (the GP-dimer), Section 6 covers GPR-accelerated NEB, and Section 7 treats GPR-accelerated minimization. Section 8 introduces the Optimal Transport GP (OT-GP) extensions that address scaling and stability. Section 9 illustrates the methods on the Muller-Brown, LEPS, and PET-MAD systems, with reproducibility details and pointers to the executable code in the supporting information.
@@ 2.3  Minimum Mode Following, the Dimer Method
We take the original formulation of the dimer for simplicity; extensions such as single-force rotation variants and Householder-based translations [75, 29] do not affect the surrogate integration developed below.
@@ 2.3.2 Translation
1:Midpoint $\mathbf{R}$, initial dimer axis $\hat{\mathbf{N}}$, separation $\Delta R$, force threshold $\epsilon_{\text{force}}$
2:repeat
3:  Place endpoints $\mathbf{R}_{1,2}=\mathbf{R}\pm\Delta R\,\hat{\mathbf{N}}$
4:  Evaluate $\mathbf{F}(\mathbf{R}_{1}),\mathbf{F}(\mathbf{R}_{2})$ on true PES $\triangleright$ 2 oracle calls
[…]
6:   $C(\hat{\mathbf{N}})\leftarrow(\mathbf{F}_{2}-\mathbf{F}_{1})\cdot\hat{\mathbf{N}}/\Delta R$ $\triangleright$ Eq. 2.6
7:   Rotate $\hat{\mathbf{N}}$ to minimize $C$ (CG, Eq. 2.7)
8:   Evaluate $\mathbf{F}(\mathbf{R}_{1})$ at new endpoint $\triangleright$ 1 oracle call per rotation
9:  until rotation converged
10:  $\triangleright$ Translation
11:  if $C(\hat{\mathbf{N}})<0$ then
12:   $\mathbf{F}^{\dagger}\leftarrow\mathbf{F}(\mathbf{R})-2[\mathbf{F}(\mathbf{R})\cdot\hat{\mathbf{N}}]\hat{\mathbf{N}} \triangleright$ Eq. 2.8, climb
13:  else
14:   $\mathbf{F}^{\dagger}\leftarrow-[\mathbf{F}(\mathbf{R})\cdot\hat{\mathbf{N}}]\hat{\mathbf{N}} \triangleright$ push toward saddle
15:  end if
16:  $\mathbf{R}\leftarrow\mathbf{R}+\text{L-BFGS step on }\mathbf{F}^{\dagger}$
17:until $|\mathbf{F}(\mathbf{R})|<\epsilon_{\text{force}}$
18:return saddle point at $\mathbf{R}$
@@ 2.4.1 The Nudged Elastic Band
A naive elastic band, where springs connect the images and the full potential force acts on each one, fails to converge to the MEP for two reasons. First, the potential force has a component along the path that drags images away from the saddle region and bunches them in low-energy basins (the "sliding-down" artifact). Second, the spring force has a component perpendicular to the path that pulls the chain off the MEP into straight-line shortcuts through high-energy regions (the "corner-cutting" artifact) [93]. The NEB eliminates both by projecting each type of force onto the subspace where it belongs. The true force acts only perpendicular to the path, driving each image toward the MEP, while the spring force acts only parallel to the path, controlling image spacing. The total NEB force on image $i$ is:
@@
This decoupling of shape optimization (perpendicular) from spacing control (parallel) is the "nudging" that gives the method its name.
@@
When the energy is monotonically increasing or decreasing through image $i$, the tangent points toward the uphill neighbor. When image $i$ sits at a local extremum (the "otherwise" case), an energy-weighted average smoothly interpolates between the two directions. The stability condition for the NEB requires that the parallel spring force be bounded by the product of the perpendicular curvature and the image spacing; the bisection tangent violates this at large $P$, while the improved tangent does not.
@@ 2.4.2 The Climbing Image and Its Connection to Minimum Mode Following
1:Initial chain $\{\mathbf{R}_{0},\ldots,\mathbf{R}_{P}\}$, spring constant $k$, CI threshold $\epsilon_{\text{CI}}$, convergence tolerance $\epsilon_{\text{tol}}$
2:CI $\leftarrow$ false
3:repeat
4:  Evaluate $V(\mathbf{R}_{i}),\mathbf{F}(\mathbf{R}_{i})$ at all movable images $\triangleright$ P1P-1 oracle calls
5:  for each image $i=1,\ldots,P-1$ do
6:   Compute tangent $\hat{\boldsymbol{\tau}}_{i}$ (Eq. 2.13)
7:   if CI is true and $i=i_{\max}$ then
8:     $\mathbf{F}_{i}\leftarrow-\nabla V+2(\nabla V\cdot\hat{\boldsymbol{\tau}}_{i})\hat{\boldsymbol{\tau}}_{i}$ $\triangleright$ Eq. 2.14
9:   else
10:     $\mathbf{F}_{i}^{\perp}\leftarrow-\nabla V+(\nabla V\cdot\hat{\boldsymbol{\tau}}_{i})\hat{\boldsymbol{\tau}}_{i}$ $\triangleright$ Eq. 2.11
11:     $F_{i}^{s}\leftarrow k(|\mathbf{R}_{i+1}-\mathbf{R}_{i}|-|\mathbf{R}_{i}-\mathbf{R}_{i-1}|)$ $\triangleright$ Eq. 2.12
12:     $\mathbf{F}_{i}\leftarrow\mathbf{F}_{i}^{\perp}+F_{i}^{s}\,\hat{\boldsymbol{\tau}}_{i}$
13:   end if
14:  end for
15:  Update all images $\mathbf{R}_{i}\leftarrow\mathbf{R}_{i}+\Delta t\,\mathbf{F}_{i}$
16:  if CI is false and $\max_{i}\|\mathbf{F}_{i}\|<\epsilon_{\text{CI}}$ then
17:   CI $\leftarrow$ true; $i_{\max}\leftarrow\arg\max_{i}V(\mathbf{R}_{i})$
18:  end if
19:until $\max_{i}\|\mathbf{F}_{i}\|<\epsilon_{\text{tol}}$
20:return converged MEP $\{\mathbf{R}_{0},\ldots,\mathbf{R}_{P}\}$